OpenStudy (anonymous):

how can i tell if the following two lines are paralell, perpendicular or neither? 6x+3y=5 and -x+2y=6

OpenStudy (lgbasallote):

good question...simplify the equations such that y has no coefficients

OpenStudy (lgbasallote):

^that is the first step

OpenStudy (anonymous):

6x + 3y = 5 3y = -6x + 5 y = -6/3x + 5/3 y = -2x + 5/3 do this for the other one.

OpenStudy (anonymous):

parallel meaning they have the same gradient, perpendicular meaning that the gradient of both lines have to times together to equal -1 or neither is which is just if their not parallel or perpendicular.

OpenStudy (anonymous):

still confused here